# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v1

@@ G.2 Theorem on Data Extraction and Proof of Corollary 2
#$\#\{\mbox{{\em BBoxER}{}}(m_{0},\omega,D_{i},a,b);i\in\{1,\dots,n\}\}\leq\sup_{D}\prod_{i=1}^{b}k_{i}.$ (A.14)
@@
Equation A.14 is an immediate consequence of Equation 1.
If $A(m_{0},D):=$BBoxER $(m_{0},\omega,D,a,b)$, then vulnerability to data extraction as in Equation 7 immediately implies that $\{E(A(m_{0},D_{i}));i\in\{1,\dots,n\}\}$ has cardinal at least $n$, and therefore $\{A(m_{0},D_{i});i\in\{1,\dots,n\}\}$ has cardinal at least $n$. If $2^{b}>n$ and $k_{i}\leq 2$, this contradicts Equation A.14.
@@ Appendix H Experimental Details
Here we provide details for the settings of the experiments in Sec. 5. The licenses for all datasets used in this work can be found in Tab. 4.
All BBO algorithms are used with their default parameters in Nevergrad [46]. At each iteration of Alg. 1, the modified model is evaluated on GSM8k train (8-shot, unless otherwise specified) and the objective function to optimize is the exact match. At the end of each run, the FinalModel is evaluated on the same objective function. Specifically, we compute the performance of $modified(m_{0},\hat{x})$ where $\hat{x}$ is the update that reaches the best objective function value during training.
@@
We apply BBoxER (Alg. 1) with rank-1 (LoRA) multiplicative parameter updates on the unembedding layer, as detailed in App. B. In both base and instruct models, the total number of parameters to optimize is $hidden\_dim+vocab\_size=4096+128,256=132,352$. Each iteration requires 2min30s on 32 A100 GPUs.
@@
    First attention layer, where we perform a broadcast multiplicative update to the Q matrix in the first attention layer by broadcasting a vector of size $4096$ to the $4096\times 4096$ Q matrix and then using the update formula as in Sec. B.1.
@@
    All attention layers, where we perform a broadcast multiplicative update on all the $Q$ matrices of all attention layers. Resulting in $n\_layers\times hidden\_dim=32\times 4096=131,072$ parameters to optimize, using the update formula as in Sec. B.2
@@
All experiments are carried out with budget $b=150$ unless otherwise specified. In addition to GSM8k test, we also evaluate the results on GSM+ (8-shot) and SVAMP (8-shot). Each iteration requires 2min30s on 32 A100 GPUs.
@@
In this experiment, D-CMA is used to directly optimize all normalization layers of Llama3.1-8B(-Instruct) and Qwen2.5-3B-Instruct (without using any low rank modification). The number of updated parameters is $n\_norm\_layers\times hidden\_dim=65\times 4096=266,240$ for Llama3.1-8B(-Instruct) and $n\_norm\_layers\times hidden\_dim=73\times 2048=149,504$ for Qwen2.5-3B-Instruct. All experiments are conducted with budget $b\in\{150,300\}$. Furthermore, we evaluate on additional mathematical reasoning benchmarks: MATH500 (4-shot), Hellaswag (0-shot), GSM+ test_mini (8-shot), ARC Easy (0-shot) and ARC Challenge (0-shot), AMC23 (0-shot). Each iteration requires 2min30s on 32 A100 GPUs for Llama3.1-8B(-Instruct) and 3min on 24 A100 GPUs for Qwen2.5-3B-Instruct.
@@ Appendix I Additional Experiments
In this section, we describe extensions of the experiments performed in Sec. 5, in particular for Bet-and-Run variants and for training on larger datasets.
@@
We have performed additional experiments on Llama3.1-8B in the spirit of Experiment 2 (Sec. 5), by modifying the output layer instead of the attention layers, using a multiplicative update by broadcasting a vector of size $vocab\_size=128,256$ as described in Sec. B.1. We also experiment with two different update constants in $\{0.01,0.0001\}$. Tab. 5 shows the results. We see improvements particularly for the larger update constant and Bet-and-Run algorithms.
[…]
@@
The theoretical results in Sec. E.4 demonstrate that Bet-and-Run algorithms enable the use of larger budgets without risks of overfitting. This has been partially validated in Experiment 2 by the results of 3-OneFifth, performing well even with greater budget (Table 1). We here extend this analysis and report Bet-and-Run results for the setup of Experiment 3 in Tab. 6. Specifically, we run BBoxER under the same setup as in Experiment 3 (detailed in App. H), and use the BetAndRun (N, b) variant of D-CMA, where we run D-CMA for N different runs with budget $b^{\prime}=N/b$, and pick the update $\hat{x}$ that performed the best across these runs. In practice, we use the runs with 5 different seeds from Tab. 2 and report the mean and standard deviation of all possible combinations $\binom{5}{N}$ with $N\in\{3,4\}$.
Table 6: Extension of Experiment 3 to Bet-and-Run algorithms, same setting as Tab. 2
@@
Llama3.1-8B-BBoxE-BetAndRun (N=3, b=450) 55.66$\pm$0.42 37.62$\pm$0.34 20.04$\pm$0.45 82.96$\pm$0.08 55.03$\pm$0.03 3.00$\pm$3.67 80.62$\pm$0.05
Llama3.1-8B-BBoxER-BetAndRun (N=3, b=900) 56.83$\pm$0.21 37.74$\pm$0.28 20.36$\pm$0.77 83.26$\pm$0.11 54.96$\pm$0.18 1.00$\pm$1.22 80.68$\pm$0.11
Llama3.1-8B-BBoxE-BetAndRun (N=4, b=600) 55.62$\pm$0.21 37.73$\pm$0.30 20.20$\pm$0.40 82.94$\pm$0.03 55.02$\pm$0.00 1.50$\pm$3.00 80.60$\pm$0.04
Llama3.1-8B-BBoxER-BetAndRun (N=4, b=1200) 56.85$\pm$0.12 37.77$\pm$0.22 20.48$\pm$0.64 83.29$\pm$0.10 54.88$\pm$0.07 0.50$\pm$1.00 80.71$\pm$0.10
@@
Llama3.1-8B-Instruct-BBoxER-BetAndRun (N=3, b=450) 78.20$\pm$0.21 55.02$\pm$0.12 36.68$\pm$0.36 78.90$\pm$0.38 55.40$\pm$0.08 23.25$\pm$2.25 79.96$\pm$0.04
Llama3.1-8B-Instruct-BBoxER-BetAndRun (N=3, b=900) 78.37$\pm$0.21 55.18$\pm$0.15 37.20$\pm$0.54 79.57$\pm$0.12 55.59$\pm$0.19 18.00$\pm$1.50 79.92$\pm$0.02
Llama3.1-8B-Instruct-BBoxER-BetAndRun (N=4, b=600) 78.26$\pm$0.03 55.03$\pm$0.10 36.52$\pm$0.24 78.77$\pm$0.34 55.42$\pm$0.07 24.00$\pm$2.00 79.96$\pm$0.03
Llama3.1-8B-Instruct-BBoxER-BetAndRun (N=4, b=1200) 78.45$\pm$0.18 55.17$\pm$0.08 37.04$\pm$0.48 79.54$\pm$0.10 55.54$\pm$0.17 17.50$\pm$0.00 79.91$\pm$0.02
@@
Qwen-2.5-3B-instruct-BBoxER-BetAndRun (N=3, b=450) 83.06$\pm$0.24 62.12$\pm$0.32 42.70$\pm$0.20 72.69$\pm$0.12 47.85$\pm$0.17 36.00$\pm$1.22 74.99$\pm$0.07
Qwen-2.5-3B-instruct-BBoxER-BetAndRun (N=3, b=900) 83.90$\pm$0.42 61.66$\pm$0.32 42.26$\pm$0.65 72.61$\pm$0.21 47.63$\pm$0.31 37.00$\pm$2.45 75.04$\pm$0.05
Qwen-2.5-3B-instruct-BBoxER-BetAndRun (N=4, b=600) 83.08$\pm$0.18 62.24$\pm$0.27 42.72$\pm$0.16 72.71$\pm$0.05 47.86$\pm$0.10 35.50$\pm$1.00 75.00$\pm$0.06
Qwen-2.5-3B-instruct-BBoxER-BetAndRun (N=4, b=1200) 84.08$\pm$0.30 61.55$\pm$0.02 42.52$\pm$0.16 72.68$\pm$0.19 47.52$\pm$0.27 36.00$\pm$2.00 75.06$\pm$0.04
Comparing the results in Tab. 6 with those in Tab. 2 (BBoxER without Bet-and-Run), we observe that in 100% runs (with or without Bet-and-Run) BBoxER improves results; in 83% of cases Bet-and-Run outperforms its base counterpart.
@@
In Tab. 7 we report the results of running BBoxER with a larger training dataset. This experiment is motivated by the theoretical prediction that the allowable budget scales linearly with the dataset size. Formally, we adopt the same experimental setup as in Experiment 3 (see App. H), but in addition to GSM8k train (7,473 samples), we also include the MATH train set (7500 samples) into the training set, thus doubling its size compared to previous experiments. We define the objective function for BBoxER to be the average of exact matches (EM) across both datasets : $metric=\frac{1}{2}EM(GSM8k)+\frac{1}{2}EM(MATH)$. We run BBoxER with a budget of up to $b=1200$ and evaluate the model every 150 iterations. As the budget increases, we observe eventual overfitting. Interestingly, we find that the Llama3.1-8B base model does not overfit on the GSM8k test set and continues to improve, contrary to its instruct variant, which begins to overfit earlier. See also Figure 3 for illustration of performance gain on some of the benchmarks.
@@
Figure 3: (full update, normalization layers) Comparison of model performance of BBoxER run with D-CMA across various benchmarks, 1 seed. See Tab. 7.
@@
Consistent with our theoretical results, our results are robust and show performance improvement with little to no overfitting even with larger budgets. We note that even when running various experiments with randomly chosen parameters, significant performance drops are rare: In Figure 1 and in Tab. 1, all data points are independent from one another (i.e. computed independently); This allows us to compute P-values (App. J) to support our claims of statistical significance of the observed performance improvements.
In order to contextualize our findings, we can also compare to post-training with GRPO which proved to be highly effective on reasoning tasks. Most works deploy GRPO on larger training sets and we were unable to find results for the models deployed in our study. However, as shown in  [141], GRPO achieved a notable 22% improvement when training a Qwen-2.5-1.5B-Instruct model on the GSM8k dataset, leading to results between 77% and 81% with 1B models, without transfer to OOD. However, such improvement jumps tend to shrink with larger models. Nevertheless, though the Qwen-2.5-3B-Instruct model we use already starts at nearly 80%, it is improved to 84% by BBoxER (Tab. 6).
There are only few works that train on GSM8k and study OOD transfer. The authors of  [85] employ an ’on-policy’ training approach, generating training samples through model prompting and subsequently fine-tuning on the resulting question-answer pairs. Their approach yields significant gains on both GSM8k and GSM+ benchmarks when evaluated in a zero-shot setting, by up to 28.8% on GSM8k with Llama3.1-8B-Instruct. However, the gains are less pronounced in the 5-shot setting, with an increase of 1.8% using self-consistency and 5 generations, figures that are close to those obtained by BBoxER on GSM8k and GSM+, as can be seen in Tab. 6.
@@
  • We computed the $p$-value of the frequency of green vs red in the columns of Tab. 1 and Tab. 5, the null hypothesis being "the probability of improvement is $\leq 0.5$".
  • For the values in Figure 1, we computed the $p$-value for the obtained frequency of successful runs (i.e., runs above baseline, vs below baseline) for budget $>20$, under the null hypothesis "the probability of improvement is $\leq 0.5$".
@@
Experiment 1: P-values in Figure 1 are $3e^{-8}$ (left), $3e^{-5}$ (right).
@@
  • Tab. 1: P-values are significant for GSM ($5e^{-3}$) and GSM+ ($0.03$), not for SVAMP.
  • Tab. 5: P-values are 0.003 for GSM, 0.5 for the transfer to GSM+, no improvement for SVAMP.

### v2

@@ G.2 Theorem on Data Extraction and Proof of Corollary˜2
$\{\mbox{{BBoxER}{}}(m_{0},\omega,D_{i},a,b);i\in\{1,\dots,n\}\}\leq\sup_{D}\prod_{i=1}^{b}k_{i}.$ (A.14)
@@
Equation˜A.14 is an immediate consequence of Equation˜1. If $A(m_{0},D):=$BBoxER $(m_{0},\omega,D,a,b)$, then vulnerability to data extraction as in Equation˜8 immediately implies that $\{E(A(m_{0},D_{i}));i\in\{1,\dots,n\}\}$ has cardinal at least $n$, and therefore $\{A(m_{0},D_{i});i\in\{1,\dots,n\}\}$ has cardinal at least $n$. If $2^{b}>n$ and $k_{i}\leq 2$, this contradicts Equation˜A.14. ∎
@@ Appendix H Experimental Details
Here we provide details for the settings of the experiments in Sec.˜5. The licenses for all datasets used in this work can be found in Tab.˜5.
[…]
@@
We apply BBoxER (Alg.˜1) with rank-1 (LoRA) multiplicative parameter updates on the unembedding layer, as detailed in App.˜B. In both base and instruct models, the total number of parameters to optimize is hidden_dim+vocab_size=4096+128,256=132,352hidden\_dim+vocab\_size=4096+128,256=132,352. Each iteration requires around 40s on 16 A100 GPUs.
@@
    First attention layer, where we perform a broadcast multiplicative update to the Q matrix in the first attention layer by broadcasting a vector of size $4096$ to the $4096\times 4096$ Q matrix and then using the update formula as in Sec.˜B.1.
@@
    All attention layers, where we perform a broadcast multiplicative update on all the $Q$ matrices of all attention layers. Resulting in $n\_layers\times hidden\_dim=32\times 4096=131,072$ parameters to optimize, using the update formula as in Sec.˜B.2
@@
All experiments are carried out with budget $b=150$ unless otherwise specified. In addition to GSM8k test, we also evaluate the results on GSM+ (8-shot) and SVAMP (8-shot). Each iteration requires around 40s on 16 A100 GPUs.
@@
In this experiment, D-CMA is used to directly optimize all normalization layers of Llama3.1-8B(-Instruct) and Qwen2.5-3B-Instruct (without using any low rank modification). The number of updated parameters is $n\_norm\_layers\times hidden\_dim=65\times 4096=266,240$ for Llama3.1-8B(-Instruct) and $n\_norm\_layers\times hidden\_dim=73\times 2048=149,504$ for Qwen2.5-3B-Instruct. All experiments are conducted with budget $b\in\{150,300\}$. Furthermore, we evaluate on additional mathematical reasoning benchmarks: MATH500 (4-shot), Hellaswag (0-shot), GSM+ test_mini (8-shot), ARC Easy (0-shot) and ARC Challenge (0-shot), AMC23 (0-shot). Each iteration requires 30s on 16 A100 GPUs for Llama3.1-8B(-Instruct) and 25s on 16 A100 GPUs for Qwen2.5-3B-Instruct.
Experiment 4
Most MIA rely on the loss (negative likelihood) of the model over the tested sample to determine membership in the training dataset [63, 64, 65]. We can formulate the membership test as follows:
$f\left(\text{NLL}\{sample\}\right)>\text{threshold}\{sample\}$
where $f$ denotes a transformation (usually continuous) applied to the sample loss. The threshold used in membership test is sample dependent, in order to account for the fact that some samples may exhibit low losses even without being part of the training set, simply because they share patterns that were seen during training.
For example, in Carlini et al. [64], the authors propose to model the distribution of model’s “confidence” with a gaussian distribution; they train shadow models to estimate the parameters $(\mu_{in},\sigma_{in},\mu_{out},\sigma_{out})$ and perform a Likelihood-ratio test:
$\frac{p(\phi(\text{NLL}\{sample\});\mu_{in},\sigma_{in})}{p(\phi(\text{NLL}\{sample\});\mu_{out},\sigma_{out})}$ $>\text{threshold}$
$\Leftrightarrow p(\phi(\text{NLL}\{sample\});\mu_{in},\sigma_{in})$ $>\text{threshold}\times p(\phi(\text{NLL}\{sample\});\mu_{out},\sigma_{out})$
Therefore, if a learning method does not impact the loss of the model over each sample, it would be difficult impossible to perform any MIA that relies on statistical tests based on the loss. We thus propose the two following metrics as proxies for the ability of membership inference attacks to succeed:
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}_{\text{inital model}}\{sample\}|$ (A.15)
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{inital model}}\{sample\}|$ (A.16)
The only distinction between Equation˜A.15 and Equation˜A.16 is which tokens are used to compute the negative log-likelihood (NLL): Equation˜A.16 computes NLL only over the (chain-of-thought, answer) pair, whereas Equation˜A.15 computes NLL over the entire input (few-shot examples, question, chain-of-thought, and answer). We aggregate the absolute differences (i.e., sum of absolute changes in sample NLL) across all samples so that positive and negative changes do not cancel each other out; although in practice the loss of the new finetuned model will be lower than the initial model. To interpret the metric: a low value indicates that the sample losses are only slightly affected by the learning method, suggesting it would be difficult to mount an attack that relies on loss changes, whereas a large value indicates that it might be possible (but not certain) to perform such an attack successfully. By showing in Table Tab.˜2 that BBoxER barely affects these metrics, and is orders of magnitude less than finetuning methods, including privacy-aware ones like DP-AdamW, we demonstrate robustness of BBoxER against a vast majority of membership inference attacks.
Below we detail all algorithms used in Tab.˜2:
  • ft: default SFT with next token prediction loss, global batch size of 128, same template used during evaluation that includes few shots examples but only the CoT and the answer contribute to the loss. AdamW is used with $(beta_{1}=0.9,beta_{2}=0.95)$ and a learning rate of $1e^{-5}$ with cosine annealing and warm-up over the first 5% steps. The training is done for 2 and 5 epochs.
  • ft-norm: similar to ft, but only the parameters of RMSNorm layers are optimized, and the learning rate is increased to $2e^{-4}$ instead.
  • DP-AdamW: differentially private training with gradient norm clipping set to $1.0$, $\delta=10^{-5}$, $L=128$, $N=7473$ samples, $292$ gradient steps (5 epochs) and a target $\epsilon=8$, the privacy accounting of [45] is used to infer the added noise scale $\sigma$. The training is done with the same template used during evaluation that includes few shots, all tokens contribute to the loss.
  • BBoxER-norm: exact setup as in experiment 3 for our algorithm, step $\delta=0.01$, only updating RMSNorm layers with an exponential multiplicative full update. Different budget values are used 150, 300, 1200. Nevergrad seed = 42.
@@ Appendix I Additional Experiments
In this section, we describe extensions of the experiments performed in Sec.˜5, in particular for Bet-and-Run variants and for training on larger datasets.
@@
We have performed additional experiments on Llama3.1-8B in the spirit of Experiment 2 (Sec.˜5), by modifying the output layer instead of the attention layers, using a multiplicative update by broadcasting a vector of size vocab_size=128,256vocab\_size=128,256 as described in Sec.˜B.1. We also experiment with two different update constants in $\{0.01,0.0001\}$. Tab.˜6 shows the results. We see improvements particularly for the larger update constant and Bet-and-Run algorithms.
Table 6: Additional Experiments. Counterpart of Tab.˜1 for the output layer: Performance difference ($\%$) after vs before BBoxER with Llama3.1-8B (Base) on GSM8k-train, evaluated on GSM8k-test (ID), and (OOD) SVAMP and GSM+, with budget $b=150$. Number of seeds in brackets. Red: detrimental runs, $1\%$. Green: beneficial runs, $1\%$. Blue rows are bet-and-run cases.
@@
The theoretical results in Sec.˜E.4 demonstrate that Bet-and-Run algorithms enable the use of larger budgets without risks of overfitting. This has been partially validated in Experiment 2 by the results of 3-OneFifth, performing well even with greater budget (Table 1). We here extend this analysis and report Bet-and-Run results for the setup of Experiment 3 in Tab.˜7. Specifically, we run BBoxER under the same setup as in Experiment 3 (detailed in App.˜H), and use the BetAndRun (N, b) variant of D-CMA, where we run D-CMA for N different runs with budget $b^{\prime}=N/b$, and pick the update $\hat{x}$ that performed the best across these runs. In practice, we use the runs with 5 different seeds from Tab.˜3 and report the mean and standard deviation of all possible combinations $\binom{5}{N}$ with $N\in\{3,4\}$.
Table 7: Extension of Experiment 3 to Bet-and-Run algorithms, same setting as Tab.˜3
@@
[…]
@@
[…]
@@
[…]
Comparing the results in Tab.˜7 with those in Tab.˜3 (BBoxER without Bet-and-Run), we observe that in 100% runs (with or without Bet-and-Run) BBoxER improves results; in 83% of cases Bet-and-Run outperforms its base counterpart.
@@
In Tab.˜8 we report the results of running BBoxER with a larger training dataset. This experiment is motivated by the theoretical prediction that the allowable budget scales linearly with the dataset size. Formally, we adopt the same experimental setup as in Experiment 3 (see App.˜H), but in addition to GSM8k train (7,473 samples), we also include the MATH train set (7500 samples) into the training set, thus doubling its size compared to previous experiments. We define the objective function for BBoxER to be the average of exact matches (EM) across both datasets : $metric=\frac{1}{2}EM(GSM8k)+\frac{1}{2}EM(MATH)$. We run BBoxER with a budget of up to $b=1200$ and evaluate the model every 150 iterations. As the budget increases, we observe eventual overfitting. Interestingly, we find that the Llama3.1-8B base model does not overfit on the GSM8k test set and continues to improve, contrary to its instruct variant, which begins to overfit earlier. See also Figure˜3 for illustration of performance gain on some of the benchmarks.
@@
Figure 3: (full update, normalization layers) Comparison of model performance of BBoxER run with D-CMA across various benchmarks, 1 seed. See Tab.˜8.
@@
Consistent with our theoretical results, our results are robust and show performance improvement with little to no overfitting even with larger budgets. We note that even when running various experiments with randomly chosen parameters, significant performance drops are rare: In Figure˜1 and in Tab.˜1, all data points are independent from one another (i.e. computed independently); This allows us to compute P-values (App.˜J) to support our claims of statistical significance of the observed performance improvements.
In order to contextualize our findings, we can also compare to post-training with GRPO which proved to be highly effective on reasoning tasks. Most works deploy GRPO on larger training sets and we were unable to find results for the models deployed in our study. However, as shown in [137], GRPO achieved a notable 22% improvement when training a Qwen-2.5-1.5B-Instruct model on the GSM8k dataset, leading to results between 77% and 81% with 1B models, without transfer to OOD. However, such improvement jumps tend to shrink with larger models. Nevertheless, though the Qwen-2.5-3B-Instruct model we use already starts at nearly 80%, it is improved to 84% by BBoxER (Tab.˜7).
There are only few works that train on GSM8k and study OOD transfer. The authors of [138] employ an ’on-policy’ training approach, generating training samples through model prompting and subsequently fine-tuning on the resulting question-answer pairs. Their approach yields significant gains on both GSM8k and GSM+ benchmarks when evaluated in a zero-shot setting, by up to 28.8% on GSM8k with Llama3.1-8B-Instruct. However, the gains are less pronounced in the 5-shot setting, with an increase of 1.8% using self-consistency and 5 generations, figures that are close to those obtained by BBoxER on GSM8k and GSM+, as can be seen in Tab.˜7.
@@
  • We computed the $p$-value of the frequency of green vs red in the columns of Tab.˜1 and Tab.˜6, the null hypothesis being "the probability of improvement is $\leq 0.5$".
  • For the values in Figure˜1, we computed the $p$-value for the obtained frequency of successful runs (i.e., runs above baseline, vs below baseline) for budget >20>20, under the null hypothesis "the probability of improvement is $\leq 0.5$".
@@
Experiment 1: P-values in Figure˜1 are $3e^{-8}$ (left), $3e^{-5}$ (right).
@@
  • Tab.˜1: P-values are significant for GSM ($5e^{-3}$) and GSM+ ($0.03$), not for SVAMP.
  • Tab.˜6: P-values are 0.003 for GSM, 0.5 for the transfer to GSM+, no improvement for SVAMP.